# Detecting Latin in Historical Books with Large Language Models: A Multimodal Benchmark

### v2

@@ Acknowledgments
This project has received funding from the European Union’s Horizon Europe programme for research and innovation under MSCA Doctoral Networks 2022, Grant Agreement No. 101120349 and Grant Agreement No. 101119511.
@@ B.2 Fuzzy Matching Algorithm in Token-level Metrics
To evaluate segment correspondence, we apply a fuzzy matching algorithm to compare lists of preprocessed ground truth tokens against predicted tokens for each sample. This approach calculates Precision, Recall, and F1-score while being robust to minor textual variations. The core matching logic is outlined in Algorithm 1.
@@
4:  $\triangleright$ Output: Precision, Recall, F1-score
@@ E.2 Prompt Experiments
Table 5: Prompt modification experiment results.
@@ E.4 Qualitative Results
Figure 7 shows an example of a mismatch caused by significant OCR noise caused by poor original image quality. Here, the post-corrected OCR of our ground truth differs so much from the OCR visual or MLLMs produced during the prediction process that not even our edit-distance based fuzzy ground truth matching can recover what is essentially a full match. This kind of error especially affects pages in the footnotes, code switching and dictionary categories, since the Latin texts in these categories tend to be printed in harder to detect fonts and layouts, which are additionally more likely to be affected by bad scan quality.

### v3

@@ Acknowledgments
We thank the Area Chair and the anonymous reviewers for their constructive comments and suggestions. This project has received funding from the European Union’s Horizon Europe programme for research and innovation under MSCA Doctoral Networks 2022, Grant Agreement No. 101120349 and Grant Agreement No. 101119511. We also acknowledge CSC – IT Center for Science, Finland, for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium.
@@ B.2 Fuzzy Matching Algorithm in Token-level Metrics
To evaluate segment correspondence, we apply a fuzzy matching algorithm to compare lists of preprocessed ground truth tokens against predicted tokens for each sample. This approach calculates Precision, Recall, and F1 score while being robust to minor textual variations. The core matching logic is outlined in Algorithm 1.
@@
4:  $\triangleright$ Output: Precision, Recall, F1 score
@@ E.2 Prompt Experiments
Table 7: Impact of prompting on Qwen2.5-VL-32B.
@@ E.4 Qualitative Results
Figure 7 shows an example of a mismatch caused by significant OCR noise caused by poor original image quality. Here, the post-corrected OCR of our ground truth differs so much from the OCR visual or MLLMs produced during the prediction process that not even our edit distance-based fuzzy ground truth matching can recover what is essentially a full match. This kind of error especially affects pages in the footnotes, code switching and dictionary categories, since the Latin texts in these categories tend to be printed in harder to detect fonts and layouts, which are additionally more likely to be affected by bad scan quality.